A car is traveling at a speed of 55 miles per hour. You want to write an algebraic expression to show how far the car will travel in a certain number of hours. What will be your constant? your variable?
Problem comes from Holt Mathematics course 1

Answer by gonzo(654)

distance = rate * time.
distance is number of miles the car has traveled.
rate = number of miles per hour the car is traveling at.
time = number of hours the car has been driven.
let D = distance.
let T = time.
let R = rate
the general equation is D = R*T
as long as the car will always be driven at 55 miles per hour, you can substitute 55 for R and that becomes your constant.
The equation then becomes D = 55 * T.
example:
you drive 5 hours.
D = 55 * T = 55 * 5 = 275 miles.
If you drive at 55 miles per hour for 5 hours, you will have driven 275 miles.